# Negative Object Presence Evaluation (NOPE) to Measure Object Hallucination in Vision-Language Models

### v1

@@ 5 Results
[…]
@@ Appendix B Reference Question Templates
No Question template
$\mathbf{NegP}$ answer
1
What is the color of the <OBJECT>?
none / nothing
2
What object is located above / under / on the left of / on the right of the <OBJECT>?
none / nothing
3
Where is the <OBJECT> located in the image?
nowhere
4
How many <OBJECT> are there in the image?
zero / 0 / none
5
Who is holding / using the <OBJECT>?
no one / nobody
6
Who is holding / using the object above / under / on the left of / on the right of the <OBJECT>?
no one / nobody
@@ Appendix F Automatic Validation Results of $\mathrm{NegP}$ VQA Data Generation
FLAN T5 XL
11
10
FLAN T5 XXL
5
17
Alpaca
44
53
FLAN Alpaca XL
25
11
ChatGLM
84
44
ChatGPT
99
98
@@ Appendix H Baselines in NOPE Benchmark
No Model References Access
Zero-shot & Few-shot
[…]
Hu et al. (2022)
https://huggingface.co/tifa-benchmark/promptcap-coco-vqa, https://huggingface.co/allenai/unifiedqa-t5-base
2 PromptCap
Hu et al. (2022)
https://huggingface.co/tifa-benchmark/promptcap-coco-vqa, https://huggingface.co/allenai/unifiedqa-t5-3b
3 BLIP-2
Li et al. (2023a)
https://huggingface.co/Salesforce/blip2-opt-2.7b
4 OpenFlamingo
Alayrac et al. (2022); Awadalla et al. (2023)
https://huggingface.co/OpenFlamingo/OpenFlamingo-9B
5 InstructBLIP
Dai et al. (2023a)
https://huggingface.co/Salesforce/instructblip-flan-t5-xl
VQA fine-tuned
1 OFA
Wang et al. (2022b)
https://huggingface.co/OFA-Sys/ofa-huge-vqa
2 BLIP
Li et al. (2022)
https://huggingface.co/Salesforce/blip-vqa-base
[…]
Li et al. (2022)
https://huggingface.co/Salesforce/blip-vqa-capfilt-large
4 ALBEF
Li et al. (2021a)
https://github.com/salesforce/ALBEF#download#Finetuned-checkpoint-for-VQA
[…]
Wang et al. (2022a)
https://huggingface.co/microsoft/git-large-vqav2
@@ Appendix I Examples of Object Hallucination in NOPE
390 Related [Uncaptioned image] Question
can you point out the location of the dog collar in the image?
GT answer
nowhere
@@
[…]
no
InstructBLIP
no
822 Related [Uncaptioned image] Question
how many menu cards have been captured in the image?
GT answer
none
@@
[…]
0
InstructBLIP
0
982 Related [Uncaptioned image] Question
where can the clouds be spotted in the image?
GT answer
nowhere
@@
9165 Partially related [Uncaptioned image] Question
who can you see using the fishing rod?
GT answer
nobody
@@
[…]
no
InstructBLIP
no one
10135 Unrelated [Uncaptioned image] Question
which color is the pillow in the image?
GT answer
nothing

### v2

@@ 5 Results
We present the results on the test set of the benchmark in Table 3. Examples of object hallucination are in Appendix I. While the VQA-finetuned baselines are slightly better at $\mathrm{NegP}$ and comparable to the zero-shot & few-shot baselines on $\mathrm{Others}$, as in Figure 7, we observe that all zero-shot and VQA-finetuned baselines notably perform much worse on $\mathrm{NegP}$ tasks that $\mathrm{Others}$ with the averaged discrepancies of $\pm$22and $\pm$18accuracy, respectively. This demonstrates that all baselines are more vulnerable and susceptible to object hallucination than standard incorrectness. In addition, less incorrectness does not entail less object hallucination. For instance, PromptCapBASE, PromptCap, and BLIP have lower scores on $\mathrm{NegP}$ than ALBEF despite outperforming it on $\mathrm{Others}$ setting. It also means that existing evaluations that solely utilize $\mathrm{Others}$ cases cannot effectively capture the models’ risk of object hallucination.
@@ 9 Ethics Statement
Acknowledgements
This work has been partially funded by PhD Fellowship Award, the Hong Kong University of Science and Technology; PF20-43679 Hong Kong PhD Fellowship Scheme, Research Grant Council, Hong Kong; and the National Research Foundation, Singapore under its AI Singapore Programme.
@@ Appendix B Reference Question Templates
No Question template $\mathbf{NegP}$ answer
1 What is the color of the <OBJECT>? none / nothing
2 What object is located above / under / on the left of / on the right of the <OBJECT>? none / nothing
3 Where is the <OBJECT> located in the image? nowhere
4 How many <OBJECT> are there in the image? zero / 0 / none
5 Who is holding / using the <OBJECT>? no one / nobody
6 Who is holding / using the object above / under / on the left of / on the right of the <OBJECT>? no one / nobody
@@ Appendix F Automatic Validation Results of $\mathrm{NegP}$ VQA Data Generation
FLAN T5 XL 11 10
FLAN T5 XXL 5 17
Alpaca 44 53
FLAN Alpaca XL 25 11
ChatGLM 84 44
ChatGPT 99 98
@@ Appendix I Examples of Object Hallucination in NOPE
390 Related [Uncaptioned image] Question can you point out the location of the dog collar in the image?
GT answer nowhere
@@
GITLARGE no
InstructBLIP no
822 Related [Uncaptioned image] Question how many menu cards have been captured in the image?
GT answer none
@@
GITLARGE 0
InstructBLIP 0
982 Related [Uncaptioned image] Question where can the clouds be spotted in the image?
GT answer nowhere
@@
9165 Partially related [Uncaptioned image] Question who can you see using the fishing rod?
GT answer nobody
@@
GITLARGE no
InstructBLIP no one
10135 Unrelated [Uncaptioned image] Question which color is the pillow in the image?
GT answer nothing